Which is a solution to the following system of equations.
3x-2y=10
4x+y=6

Answer by mananth(16946)

3 x + -2 y = 10 .............1
4 x + 1 y = 6 .............2

multiply (1)by 1
Multiply (2) by 2
3 x -2 y = 10
8 x + 2 y = 12

Add the two equations

11 x = 22

/ 11

x=2

plug value of x in (1)

3 x -2y = 10
6-2 y = 10
-2 y = 10 + -6
-2 y = 4
y = -2